# Constraining the Swift Memory Burden Effect with GW250114-like Events

### v1

@@ Abstract
Black hole spectroscopy allows to infer the properties of the remnant of a binary black hole coalescence. Motivated by the recent proposal that a black hole’s information load can alter its classical response to small perturbations, an effect known as the swift memory burden, we develop a minimal phenomenological framework to analyze the ringdown of a binary black hole merger and confront it with the data from the GW250114 event. We perform a Bayesian analysis combining the frequencies of the (220) and (440) quasi-normal modes and obtain a lower bound $\log_{10}p\gtrsim 2$, where $p$ controls how the gaps reopen when the black hole’s master mode occupation departs from the critical value. Moreover, using a Fisher information matrix (high signal-to-noise ratio) approximation, we forecast the lower bound $\log_{10}p\gtrsim 5$ for a GW250114-like event observed with Cosmic Explorer. Our results disfavour rapid gap reopening, shedding light on how the swift memory burden effect can be probed with current and next-generation detectors.
@@ I Introduction
Guided by the SMB picture, our analysis proceeds in four steps. First, we provide a brief review of the SMB in Section II. In Section III, we develop a minimal single-mode ringdown model that includes the SMB-induced suppression and the QNM frequency redshift due to the SMB effect. Then in Section IV, we confront this parameterization by performing a Bayesian parameter estimation using the QNM data from the GW250114 event [11]. Next, we assess the prospects for next-generation detectors by computing Fisher-matrix forecasts [57] for a GW250114-like source observed by Cosmic Explorer (CE) [58, 59] in Section V. Finally, we discuss and summarize in Section VI.
@@
$S(f;\mu,p)=\begin{cases}\displaystyle\Bigl[\mu\Bigl(1-\frac{f}{f_{R}}\Bigr)\Bigr]^{\frac{1}{p-1}},&f\leq f_{R},\\[6.0pt] 0,&f\geq f_{R}.\end{cases}$ (1)
@@ III Peak shift due to SMB
Notice that the suppression factor is around unity for $p\gg 1$ if the BH information load is full ($\mu\ll 1$). For $p\sim\mathcal{O}(1)$, the suppression is extremely small and almost no radiation will be emitted. A physical choice hence corresponds to $p\gg 1$, in this case, the peak of the ringdown spectrum will be redshifted due to the SMB suppression factor such that
@@ IV Bayesian inference based on GW250114
Figure 1: Left panel: Posterior constraints on the SMB model parameters using the GW250114 event data (blue and orange), together with a Fisher-matrix forecast for a GW250114-like source observed by Cosmic Explorer. Blue: LVK posterior for GW250114 combining both the 220 and 440 modes data; orange: including only the 220 mode data. Right panel: Probability distribution of the QNM frequency deviation parameter $\delta f$, for the 220 mode, obtained from the Fisher-matrix forecast for a GW250114-like source observed by CE. The red dashed line marks the mean value.
@@ V Fisher Matrix Forecasts for GW250114-like Events
To estimate the expected precision with which future ground-based detectors such as the Cosmic Explorer (CE) could measure potential deviations from the GR QNM frequencies, we performed a Fisher-matrix forecast [57] based on a GW250114-like event.
@@
where $(a|b)$ denotes the standard noise-weighted inner product over frequency, and $\theta_{i}$ are the model parameters. The signal-to-noise ratio (SNR) and Fisher matrix were computed for a CE detector configuration. The corresponding inverse Fisher matrix provides the covariance matrix $\Sigma=\Gamma^{-1}$, from which we derived the $1\sigma$ uncertainties.
To visualize the expected parameter correlations, we sampled $5\times 10^{5}$ points from the multivariate Gaussian defined by $\Sigma$ and the fiducial mean values. The right panel of Fig. 1 shows the marginalized distribution of $\delta f$. The analysis indicates that, for a GW250114-like event observed by CE, the deviation parameter $\delta f$ can be constrained to the order of $\sigma_{\delta f}\sim 10^{-4}$, depending on the assumed network sensitivity. This result highlights the capability of next-generation detectors to perform precision tests of the deviation of the QNM frequencies with unprecedented accuracy.
Again, we perform another Bayesian inference based on the Fisher-matrix forecast result, as shown in the right panel of Fig. 1 for $\delta f$ and also the corresponding bound of $p$ in the left panel. It can be seen that the CE Fisher forecast (green) results of $p$ move to higher values, indicating a prospective bound $\log_{10}p\gtrsim 5$, three orders of magnitude better than the LVK result. Because the SMB suppression factor scales as $(1-f/f_{R})\mu]^{1/(p-1)}$, these bounds force the power-law index 1/(p1)1/(p-1) to be very small. Since the observed frequency shift of GW250114-like events is small, it implies $\mathcal{O}(1)$ suppression near $f_{R}$.
@@ VI Summary and discussion
We have proposed and tested a minimal, data-driven framework to probe the SMB with the ringdown signal of binary BH merger. In the SMB picture, assisted gaplessness makes a large family of memory modes nearly degenerate when a merger drives the system. The information load resists depletion into radiation, pushing the effective resonance toward the infrared and narrowing the resonance window. We encoded this physics in a single-mode ringdown model in which the spectral peak is slightly redshifted while the overall intensity near the unburdened resonance is suppressed. Using LVK posteriors for GW250114 and combining the (220) and (440) modes, our Bayesian analysis yields a lower bound $\log_{10}p\gtrsim 2$. On the other hand, Fisher-matrix forecasts for a CE observation of a GW250114-like source indicate prospective constraints around $\log_{10}p\gtrsim 5$. Together, these results show that current data already disfavor strongly burdened responses with rapidly reopening gaps, and that next-generation sensitivity will further constrain the parameters.

### v2

@@ Abstract
Black hole spectroscopy allows to infer the properties of the remnant of a binary black hole coalescence. Motivated by the recent proposal that a black hole’s information load can alter its classical response to small perturbations, an effect known as the swift memory burden, we develop a minimal phenomenological framework to analyze the ringdown of a binary black hole merger and confront it with the data from the GW250114 event. We perform a Bayesian analysis combining the frequencies of the (220) and (440) quasi-normal modes and obtain a lower bound $\log_{10}p\gtrsim 2$, where $p$ controls how the gaps reopen when the black hole’s master mode occupation departs from the critical value. Moreover, using a Fisher information matrix (high signal-to-noise ratio) approximation, we forecast the lower bound $\log_{10}p\gtrsim 3$ for a GW250114-like event observed with Cosmic Explorer or Einstein Telescope. Our results disfavour rapid gap reopening, shedding light on how the swift memory burden effect can be probed with current and next-generation detectors.
@@ I Introduction
Guided by the SMB picture, our analysis proceeds in four steps. First, we provide a brief review of the SMB in Section II. In Section III, we develop a minimal single-mode ringdown model that includes the SMB-induced suppression and the QNM frequency redshift due to the SMB effect. Then in Section IV, we confront this parameterization by performing a Bayesian parameter estimation using the QNM data from the GW250114 event [11]. Next, we assess the prospects for next-generation detectors by computing Fisher-matrix forecasts [57] for a GW250114-like source observed by Cosmic Explorer (CE) [58, 59] or Einstein Telescope (ET) [60, 61] in Section V. Finally, we discuss and summarize in Section VI.
@@
$S(f;\mu,p)=\begin{cases}\displaystyle\Bigl[\mu\Bigl(1-\frac{f}{f_{R}}\Bigr)\Bigr]^{\frac{1}{p-1}},&f\leq f_{R},\\[6.0pt] 1,&f\geq f_{R}.\end{cases}$ (1)
@@ III Peak shift due to SMB
Notice that the suppression factor is around unity for $p\gg 1$ if the BH information load is full ($\mu\ll 1$). For $p\sim\mathcal{O}(1)$, the suppression factor is extremely small and almost no radiation will be emitted. A physical choice hence corresponds to $p\gg 1$, in this case, the peak of the ringdown spectrum will be redshifted due to the SMB suppression factor such that
@@ IV Bayesian inference based on GW250114
Figure 1: Left panel: Posterior constraints on the SMB model parameters using the GW250114 event data (blue and orange), together with a Fisher-matrix forecast for a GW250114-like source observed by Cosmic Explorer. Blue: LVK posterior for GW250114 combining both the 220 and 440 modes data; orange: including only the 220 mode data. Right panel: Probability distribution of the QNM frequency deviation parameter $\delta f$, for the 220 mode, obtained from the Fisher-matrix forecast for a GW250114-like source observed by CE/ET. The red dashed line marks the mean value.
@@ V Fisher Matrix Forecasts for GW250114-like Events
To estimate the expected precision with which future ground-based detectors such as CE and ET could measure potential deviations from the GR QNM frequencies, we performed a Fisher-matrix forecast [57] based on a GW250114-like event.
@@
where $(a|b)$ denotes the standard noise-weighted inner product over frequency, and $\theta_{i}$ are the model parameters. The signal-to-noise ratio (SNR) and Fisher matrix were computed for CE/ET detector configurations. The corresponding inverse Fisher matrix provides the covariance matrix $\Sigma=\Gamma^{-1}$, from which we derived the $1\sigma$ uncertainties.
To visualize the expected parameter correlations, we sampled $5\times 10^{5}$ points from the multivariate Gaussian defined by $\Sigma$ and the fiducial mean values. The right panel of Fig. 1 shows the marginalized distribution of $\delta f$. The analysis indicates that, for a GW250114-like event observed by CE/ET, the deviation parameter $\delta f$ can be constrained to a few thousandth, depending on the assumed network sensitivity. This result highlights the capability of next-generation detectors to perform precision tests of the deviation of the QNM frequencies with unprecedented accuracy.
Again, we perform another Bayesian inference based on the Fisher-matrix forecast result, as shown in the right panel of Fig. 1 for $\delta f$ and also the corresponding bound of $p$ in the left panel. It can be seen that the CE/ET Fisher forecast (red/green) results of $p$ move to higher values, indicating a prospective bound $\log_{10}p\gtrsim 3$, one order of magnitude better than the LVK result. Because the SMB suppression factor scales as $(1-f/f_{R})\mu]^{1/(p-1)}$, these bounds force the power-law index 1/(p1)1/(p-1) to be very small. Since the observed frequency shift of GW250114-like events is small, it implies $\mathcal{O}(1)$ suppression near $f_{R}$.
@@ VI Summary and discussion
We have proposed and tested a minimal, data-driven framework to probe the SMB with the ringdown signal of binary BH merger. In the SMB picture, assisted gaplessness makes a large family of memory modes nearly degenerate when a merger drives the system. The information load resists depletion into radiation, pushing the effective resonance toward the infrared and narrowing the resonance window. We encoded this physics in a single-mode ringdown model in which the spectral peak is slightly redshifted while the overall intensity near the unburdened resonance is suppressed. Using LVK posteriors for GW250114 and combining the (220) and (440) modes, our Bayesian analysis yields a lower bound $\log_{10}p\gtrsim 2$. On the other hand, Fisher-matrix forecasts for a CE/ET observation of a GW250114-like source indicate prospective constraints around $\log_{10}p\gtrsim 3$. Together, these results show that current data already disfavor strongly burdened responses with rapidly reopening gaps, and that next-generation sensitivity will further constrain the parameters.